# DiffStitch: Boosting Offline Reinforcement Learning with Diffusion-based Trajectory Stitching

### v1

@@ 4.1.1 Step Estimation Module
[…]
@@ 4.1.2 State Stitching Module
$\tau_{s,m}=(\bm{s}_{T},\underbrace{\texttt{[MASK]},\ldots,\texttt{[MASK]}}_{\Delta\text{ masks}},\bm{s}^{\prime}_{1},.,\bm{s}^{\prime}_{H-1-\Delta}),$ (5)
@@
$=(\bm{s}_{T},\underbrace{\widetilde{s_{1}},\ldots,\widetilde{\bm{s}_{\Delta}}}_{\Delta\text{ stitching states}},\bm{s}^{\prime}_{1},.,\bm{s}^{\prime}_{H-1-\Delta}),$
@@ 5.1 Experiment Settings
For the generative model, the horizon of generation is set to 100 for MuJoCo tasks, and 55 for Adroit-pen tasks. The denoising step $K$ is set to 100. We implement the inverse dynamics model $f_{\psi}$ with a 2-layer MLP, and the dynamic model $f_{\omega}$, reward model $f_{\phi}$ with a 4-layer MLP. For the stitching, we set the qualification threshold used for data selection $\delta$ ranging from [1,16][1,16][ 1 , 16 ] depending on the dataset. Our model is trained on a device with A5000 GPUs(24GB GPU memory, 27.8 TFLOPS computing capabilities, AMD EPYC 7371 16-Core Processor, optimized by Adam (Kingma & Ba, 2014) optimizer.
@@ 5.5 Parameter Study
DiffStitch enhances the effectiveness of offline RL algorithms by generating augmented data. In the training of offline RL algorithms, both the augmented dataset and original offline dataset are used. Consequently, the ratio of the original data to the augmented data might impact the performance of DiffStitch. To investigate that, we conduct experiments on Halfcheetah-Medium-expert with different ratios. Specifically, we set the ratio to $0\}$ : 1 , 1 : 1 , 2 : 1 , 4 : 1 , 9 : 1 , 1 : 0 }, and other settings remain the same to test the performance. The results are illustrated in Table 2. Except for the effectiveness of the augmentation, we can also observe in Table 2 that both an excessively large ratio and an excessively small ratio have detrimental effects on the performance of DiffStitch. We assume the reasons are as follows: When ratio is excessively large(1:0), the benefit brought by augmentation is abandoned. When data ratio is excessively small(0:1), too much noise are introduced and agents cannot effectively learn a policy.

### v2

@@ 4.1.1 Step Estimation Module
where $\Delta$ denotes the number of time steps to connect $\bm{s}_{T}$ and $\bm{s}^{\prime}_{1}$. Obviously, $\Delta\leq H-2$ Here, we implement $\text{sim}(\cdot)$ with cosine similarity.
@@ 4.1.2 State Stitching Module
$\tau_{s,m}=(\bm{s}_{T},\underbrace{\texttt{[MASK]},\ldots,\texttt{[MASK]}}_{\Delta\text{ masks}},\bm{s}^{\prime}_{1},...,\bm{s}^{\prime}_{H-1-\Delta}),$ (5)
@@
$=(\bm{s}_{T},\underbrace{\widetilde{s_{1}},\ldots,\widetilde{\bm{s}_{\Delta}}}_{\Delta\text{ stitching states}},\bm{s}^{\prime}_{1},...,\bm{s}^{\prime}_{H-1-\Delta}),$
@@ 5.1 Experiment Settings
For the generative model, the horizon of generation is set to 100 for MuJoCo tasks, and 55 for Adroit-pen tasks. The denoising step $K$ is set to 100. We implement the inverse dynamics model $f_{\psi}$ with a 2-layer MLP, and the dynamic model $f_{\omega}$, reward model $f_{\phi}$ with a 4-layer MLP. For the stitching, we set the qualification threshold used for data selection $\delta$ ranging from $[1,16]$ depending on the dataset. Our model is trained on a device with A5000 GPUs(24GB GPU memory, 27.8 TFLOPS computing capabilities, AMD EPYC 7371 16-Core Processor, optimized by Adam (Kingma & Ba, 2014) optimizer.
@@ 5.5 Parameter Study
DiffStitch enhances the effectiveness of offline RL algorithms by generating augmented data. In the training of offline RL algorithms, both the augmented dataset and original offline dataset are used. Consequently, the ratio of the original data to the augmented data might impact the performance of DiffStitch. To investigate that, we conduct experiments on Halfcheetah-Medium-expert with different ratios. Specifically, we set the ratio to $\{0:1,1:1,2:1,4:1,9:1,1:0\}$, and other settings remain the same to test the performance. The results are illustrated in Table 2. Except for the effectiveness of the augmentation, we can also observe in Table 2 that both an excessively large ratio and an excessively small ratio have detrimental effects on the performance of DiffStitch. We assume the reasons are as follows: When ratio is excessively large(1:0), the benefit brought by augmentation is abandoned. When data ratio is excessively small(0:1), too much noise are introduced and agents cannot effectively learn a policy.